A shell acquires the initial velocity v = 320 m/s, having made n = 2.0 turns inside the barrel whose length is equal to l = 2.0 m. Assuming that the shell moves inside the barrel with a uniform acceleration, find the angular velocity of its axial rotation at the moment when the shell escapes the barrel.
1
Expert's answer
2020-10-13T07:13:36-0400

Solution

In this case angular acceleration can be expressed as below

"\omega=\frac{2\pi nv}{l}"

Putting values

n=2

v=320m/s

l=2

"\omega=\frac{2\times 3.14\times2\times320}{2}"

"\omega=2009.6" "rad\/s^2"